# A Comprehensive Survey on Network Traffic Synthesis: From Statistical Models to Deep Learning

### v1

@@ IV-B Weibull distribution
As discussed earlier, the limitations of Poisson processes for modeling network traffic led to exploring other distributions. One such promising model was the Weibull distribution, mainly due to its flexibility to model both heavy and non-heavy tailed distributions [11]. The Weibull distribution is a continuous probability distribution where the probability density function is as follows:
@@
Here, $x\geq\mu$ and $\gamma,\alpha>0$ where $\gamma$ is the shape parameter, $\mu$ is the location parameter and $\alpha$ is the scale parameter. The case where $\mu=0$ and $\alpha=1$ is called the standard Weibull distribution, while the case where $\mu=0$ is called the 2-parameter Weibull distribution.
@@
Initial studies focused on the application of Weibull distributions to wireless TCP/IP traffic. Kullback et al. [79], highlighted the inadequacy of conventional distributions in modeling the marginal distribution of wireless TCP flow inter-arrival times. The authors proposed a piecewise Weibull distribution and a canonical multinomial cascade model to capture the salient characteristics of wireless TCP flow inter-arrivals. This model achieved an acceptable accuracy for predicting packet loss rates. However, it states that while wireless TCP/IP can be partially modeled, particularly the inter-arrival times of new flows and the number of packets per flow, the arrival process of packets within a flow is not tractable due to correlations, periodicity, and inconsistency.
@@
Further research was done in modeling internet traffic across various network tiers utilizing the Weibull distribution. Arfeen et al [12] focused on how the two-parameter Weibull distribution can be used in internet traffic modeling. It demonstrates how the Weibull distribution captures the transformation of inter-arrival process of all structural components of internet traffic (packets, flows, and sessions) as traffic moves from access to the core network. Furthermore, the authors show that given a suitable scale parameter (media or tier specific) the Weibull shape parameter can be used to zoom in from session to flow and to packet-level inter-arrivals.
@@
Later in 2019, A. Arfeen et al. [11] provide the analytical justification for why the traffic converges to Weibull distribution as traffic moves from access to core links. This convergence is attributed to the Weibull distribution’s flexibility in capturing the stochastic properties of internet traffic at different levels (packets, flows, sessions). The study presents a comprehensive analysis of all structural components of internet traffic, linking fractal renewal processes to the modeling of correlated traffic. Additionally, the research validates Weibull-based count and interval traffic data modeling using both data fitness and queuing analysis.
IV-C Markov Model
This is a mathematical framework for modeling systems that undergo transitions from one state to another in a probabilistic manner based on the Markov property. Markov property states that the future state of the system depends only on the current state and not on the sequence of events that preceded it.
@@ IV-C Markov Model
In contrast to the previous models, Markov models offer a complementary approach, particularly capturing the temporal dependencies and oscillatory behavior often observed in network traffic. Despite the inability to achieve true LRD or self-similarity, Markov models can be used to model internet traffic due to the possibility of exploiting analytical techniques to predict network performance, which according to [105] is important as it is the ultimate goal when adopting models to either study existing networks or design new ones.
@@
Most work that proposed various traffic models to capture LRD prior to the work of L. Muscariell et al. [105] lacked practical application in network design, planning, and performance analysis has been limited due to the complex mathematical structures underlying these stochastic processes. The authors further highlight that long-term correlations beyond a certain threshold may not significantly impact system performance and hence suggest simpler models like the Markov Modulated Poisson Process (MMPP) can be effectively employed in many cases. Accordingly, the authors introduce a 5-parameter hierarchical MMPP to model real-world internet traffic in multiple levels, including sessions, flows, and packets. Three of these parameters can be directly mapped onto average traffic parameters, such as the average flow arrival rate, the average number of packets per flow, and the average arrival rate of packets within flows. The other two parameters, the average number of flows per session, and the arrival rate of new sessions define the notion of session and are used to control the Hurst parameter of the synthetic traffic on the considered scaling range. The authors conclude that the behavior of this synthetic traffic in a queue with either finite or infinite buffers is coherent with real-world internet traffic.
@@
In 2012, Y. Xie et al.  [163] proposed a novel approach to model the oscillatory behavior of network traffic using Nested Hidden Markov Models (NHMMs) with variable state-duration. The proposed model synthesizes traffic traces that match the input trace’s characteristics such as temporal structure of oscillating behavior, statistical properties, self-similarity, queuing behavior, and multiscale energy distribution. The authors address the limitations of traditional Hidden Markov Models (HMM) and MMPPs, which can suffer from overfitting and increased computational complexity when increasing the number of states. NHMMs, on the other hand, provide a more effective and flexible approach. The model employs a hierarchical structure, consisting of two nested hidden Markov chains and an observable process. The first-layer hidden Markov chain, with variable state-duration, controls the overall oscillatory process. The second-layer hidden Markov chain models the local fluctuations within the oscillatory process, conditioned on the first-layer Markov chain. The results demonstrate the superior performance of NHMMs with variable state-duration in capturing the time structure of real network traffic. Compared to empirical statistical methods, multifractal wavelet methods, HMMs, and MMPPs, NHMMs more accurately match the oscillation characteristics, self-similarity, and queuing behavior of the input traces.
@@
The authors in [43] utilize hidden Markov models to model inter packet time and packet size for Internet traffic from multiple application-layer protocols. The authors evaluate the model’s ability to learn, generate, and predict realistic packet-level traffic patterns through automated analysis of empirical traffic traces, focusing on the marginal distributions, auto-covariance, and cross-covariance of inter-packet time and packet size. The work shows that the proposed model is capable of predicting the short-term future behavior of the analyzed traffic based on limited monitored traffic. Unlike previous work, which primarily focused on HTTP, this study extends the application of HMMs to other application-layer protocols such as SMTP.
Among the work that presented the use of Markov chains to synthesize network traffic for practical applications,  [50] explore the application of Markov chains for detecting malicious network behavior. This analyzed the network traffic behavior between several botnets and normal network traffic and designed a behavioral network only considering the size, duration, and periodicity of each flow. By creating a dataset using virtual machines that ran real botnets, the work recognizes that the most important behaviors to be detected on a botnet are those of Command and Control (C&C) channels. All the flows of the same C&C channel are aggregated by representing a connection of flows that share a 4-tuple of source IP address, destination IP address, destination port, and protocol. The behavioral model then identifies several states for the C&C channels based on the size, duration, and periodicity for each flow and represents the transitions between states as a chain. A concatenated chain of states is then used to generate a Markov chain for the C&C channel, which is then stored on disk as a representation of that behavior. Additionally, the probability of generating the original chain of states using its own Markov chain is also stored for the purpose of botnet detection. The detection is done by finding out whether a new chain of states was generated by the Markov chain of a pre-trained model. The work recognizes that traffic from a C&C channel in one botnet will be different to other botnets. Furthermore, part of the testing chains might be very similar while other parts will be completely different. The authors designed a detection method that compares a chain of states from a 4-tuple with a Markov chain that does not have all of its states. This method obtained an F-Measure up to 93%.
IV-D Autoregression
@@
An autoregressive (AR) model is a type of time series model where the current value of a time series is expressed as a linear combination of its previous values (lags) and a random error term. An autoregressive model of order $p$, denoted as AR(p), is defined as:
@@
where, $X_{t}$ is the value of the time series at time $t$, $c$ is a constant term, $\phi_{1},\phi_{2},\ldots,\phi_{p}$ are the autoregressive coefficients, $p$ is the number of previous time steps (lags) considered in the model, and $\epsilon_{t}$ is a random error term (white noise) with mean 00 and constant variance $\sigma^{2}$. These models are widely used in time series analysis and forecasting.
ARIMA (AutoRegressive Integrated Moving Average) models extend the autoregressive model by considering the moving average and non-stationarity. FARIMA (Fractional ARIMA) extends ARIMA by introducing fractional differencing to account for long-memory dependence.
Several later works further explored ARIMA/FARIMA models and proposed improvements. Shu et al. [135] investigates the use of FARIMA models to accurately predict network traffic. The study acknowledges that for an accurate prediction, the model should be able to capture both Short Range Dependence (SRD) and Long Range Dependence (LRD) characteristics seen in network traffic. The authors state that the FARIMA models, the Markov modulated processes, TES models, and scene-based models have the capability to capture both SRD and LRD. However, the authors argue that the complexity of FARIMA model fitting has hindered their application for prediction. To address this, the authors propose a simplified FARIMA model fitting procedure. They observe that the order (p, q) of the fitted FARIMA model often remains relatively stable over time, even for extended periods. This allows for a streamlined fitting process. Additionally, the study introduces an adapted prediction method that provides an upper probability limit for the forecast. Their results highlight the suitability of FARIMA models over AR models in h-step forecasting. The study concludes that the FARIMA models offer better accuracy while requiring fewer parameters.
ARIMA/GARCH [179], a non-linear time series model that combines the linear ARIMA with conditional Generalized Auto Regressive Conditional Heteroscedasticity (GARCH) to model and predict network traffic. The authors present that an accurate network traffic prediction model should have the ability to capture SRD and LRD, Self-similarity on a large-time scale, and multifractal on a small-time scale present in network traffic. The authors identify that the exponential decay of the autocorrelation function of ARIMA captures only the SRD characteristic, but fails to capture LRD characteristics. Similarly, FARIMA models can capture both SRD and LRD but fail to capture the multifractal characteristics. Multifractal Wavelet Model (MWM), which captures multifractal characteristics, fails to predict network traffic. Furthermore, network traffic also exhibits non-stationary and non-linear properties. Addressing all the issues with previously explored models, ARIMA/GARCH model was introduced. The proposed GARCH component contributes to the dynamic variance, where variance varies over time. This conditional variance property captures the network’s bursty characteristics. Furthermore, the study presents a new procedure to estimate the parameters of the ARIMA/GARCH model and an adaptive prediction scheme. The prediction performance shows that this model outperforms the traditional FARIMA model.
IV-E Fractional Brownian Motion
A fractional Brownian motion (FBM) is defined as a centered Gaussian process $B_{H}(t)$, $t\geq 0$ with the covariance function,
@@ IV-E Fractional Brownian Motion
Fractional Brownian Motion (FBM) has been considered a suitable choice for modeling network traffic due to its inherent self-similarity, a property that mirrors modern network traffic. The authors in [116] study the accuracy of simulation for network traffic using FBM. The authors describe simulation methods for FBM based on its spectral representation and the property of stationary increments. The indicated methods were used for the simulation of self-similar traffic and the loading process of telecommunication networks. The simulated FBM is then used as an input process to a linear system. The goal is to construct a model that accurately approximates FBM while also considering the system’s response. While the Gaussian nature of FBM aids in analyzing queuing behavior, its limitations in capturing short-term correlations and rich scaling properties of real traffic have led researchers to explore multifractal models [105]. However, these models were also difficult to manage due to their analytical complexity.
IV-F Other
IV-F1 Hierarchical State-Machine-Based Traffic Models
These have been introduced as a more accurate approach to modeling control-plane traffic in cellular networks, particularly for 4G LTE and 5G. Traditional models, such as Poisson processes, have proven inadequate due to their inability to capture the bursty nature and long-tailed distributions observed in control-plane traffic.
@@ IV-F1 Hierarchical State-Machine-Based Traffic Models
The feasibility of addressing these limitations by employing a two-level hierarchical state-machine structure to accurately model control-plane traffic was presented by J. Meng et al. [98]. The authors show that this structure enables the capture of event dependencies among different types of control-plane events. Furthermore, by using a Semi-Markov process, the model can accurately represent the time spent by a User Equipment in each state, overcoming the limitations of the Markov model. Additionally, an adaptive clustering scheme is included to account for variations in traffic behavior based on factors such as device type and time of day. The proposed model is experimentally validated by comparing the synthetic traffic generated by the model against real-world traffic traces.
IV-F2 K-Nearest Neighbors

### v2

@@ III-B Weibull distribution (WD)
As discussed earlier, the limitations of PPs for modeling network traffic led to exploring other distributions. Overcoming the limitations of PPs, WD model was proposed mainly due to its flexibility to model both heavy and non-heavy tailed distributions [11]. The WD is a continuous probability distribution where the probability density function is as follows:
@@
Here, $x\geq\mu$ and $\gamma,\alpha>0$ where $\gamma$ is the shape parameter, $\mu$ is the location parameter and $\alpha$ is the scale parameter. The case where $\mu=0$ and $\alpha=1$ is called the standard WD, while the case where $\mu=0$ is called the 2-parameter WD. Initial studies focused on the application of WDs to wireless TCP/IP traffic. Kullback et al. [93], highlighted the inadequacy of conventional distributions in modeling the marginal distribution of wireless TCP flow inter-arrival times. The authors proposed a piecewise WD and a canonical multinomial cascade model to capture the salient characteristics of wireless TCP flow inter-arrivals. This model achieved an acceptable accuracy for predicting packet loss rates. However, it states that while wireless TCP/IP can be partially modeled, particularly the inter-arrival times of new flows and the number of packets per flow, the arrival process of packets within a flow is not tractable due to correlations, periodicity, and inconsistency.
@@
Further research was done in modeling internet traffic across various network tiers utilizing the WD. Arfeen et al [12] focused on how the two-parameter WD can be used in internet traffic modeling. It demonstrates how the WD captures the transformation of inter-arrival process of all structural components of internet traffic (packets, flows, and sessions) as traffic moves from access to the core network. Furthermore, the authors show that given a suitable scale parameter (media or tier specific) the Weibull shape parameter can be used to zoom in from session to flow and to packet-level inter-arrivals.
@@
Later in 2019, A. Arfeen et al. [11] provide the analytical justification for why the traffic converges to WD as traffic moves from access to core links. This convergence is attributed to the WD’s flexibility in capturing the stochastic properties of internet traffic at different levels (packets, flows, sessions). The study presents a comprehensive analysis of all structural components of internet traffic, linking fractal renewal processes to the modeling of correlated traffic. Additionally, the research validates Weibull-based count and interval traffic data modeling using both data fitness and queuing analysis.
In [137], the authors propose a fingerprinting-based indoor positioning method that models bluetooth signal using a WD to address the limitations of sparse Received Signal Strength Indicator (RSSI) during the training phase. By estimating the shape and scale parameters, the approach captures the full RSSI domain more effectively than conventional occurrence-based methods. For the positioning phase, a histogram maximum likelihood estimator based on bayesian theory is employed to improve localization accuracy and reliability.
Limitations: Although Weibull-based models significantly improve marginal distribution fitting compared to Poisson models, their limitations emerge when moving beyond distributional accuracy. As shown in [93], WDs struggle to represent packet arrivals within flows due to correlation and periodicity effects. Moreover, while [12] demonstrates that Weibull parameters can link traffic behavior across network tiers, this linkage depends on tier-specific scale parameters, reducing model universality. These observations suggest that WDs are descriptive rather than generative in a temporal sense, capturing what traffic “looks like” statistically but not how it evolves dynamically.
Summary: Overall, WDs occupy an important middle ground in statistical traffic modeling. The works discussed demonstrate their ability to flexibly capture inter-arrival characteristics across packets, flows, and sessions and to explain how traffic distributions transform from access to core networks [12, 11]. While WDs do not explicitly model temporal dependence, they provide strong empirical grounding for understanding traffic structure and serve as effective components for hybrid or hierarchical modeling approaches.
III-C Markov Model(MM)
@@
This is a mathematical framework for modeling systems that undergo transitions from one state to another in a probabilistic manner based on the Markov property. Markov property states that the future state of the system depends only on the current state and not on the sequence of events that preceded it. In contrast to the previous models, MMs offer a complementary approach, particularly capturing the temporal dependencies and oscillatory behavior often observed in network traffic. Despite the inability to achieve true LRD or self-similarity, MMs can be used to model internet traffic due to the possibility of exploiting analytical techniques to predict network performance, which according to [121] is important as it is the ultimate goal when adopting models to either study existing networks or design new ones.
@@
Most traffic models proposed to capture LRD before Muscariello et al. [121] had limited practical use in network design due to complex mathematical structures. The authors argue that long-term correlations beyond a threshold may not significantly affect performance, suggesting simpler models like Markov-Modulated PP (MMPP) are often sufficient. They introduce a 5-parameter hierarchical MMPP to model traffic across sessions, flows, and packets. Three parameters map directly to average traffic metrics (flow arrival rate, packets per flow, and packet arrival rate), while the other two (flows per session and session arrival rate) define sessions and control the Hurst parameter. The resulting synthetic traffic shows queueing behavior consistent with real-world internet traffic.
@@
In 2012, Y.Xie et al. [185] proposed Nested Hidden MMs (NHMMs) with variable state-duration to model oscillatory network traffic. The model synthesizes traces that preserve temporal oscillations, self-similarity, statistical properties, queuing behavior, and multiscale energy distribution. Unlike Hidden Markov Models (HMMs) and MMPPs, which face overfitting and high complexity with more states, NHMMs use a hierarchical structure of two nested chains: the first controls overall oscillations, and the second models local fluctuations. Results show NHMMs outperform empirical, multifractal, HMM, and MMPP methods in capturing traffic dynamics.
@@
The authors in [47] utilize HMMs to model inter packet time and packet size for Internet traffic from multiple application-layer protocols. The authors evaluate the model’s ability to learn, generate, and predict realistic packet-level traffic patterns through automated analysis of empirical traffic traces, focusing on the marginal distributions, auto-covariance, and cross-covariance of inter-packet time and packet size. The work shows that the proposed model is capable of predicting the short-term future behavior of the analyzed traffic based on limited monitored traffic. Unlike previous work, which primarily focused on HTTP, this study extends the application of HMMs to other application-layer protocols such as Simple Mail Transfer Protocol (SMTP).
Garcia et al. [59] applied Markov chains to detect malicious traffic by modeling botnet Command and Control (C&C) behavior. Using flows defined by size, duration, and periodicity, they built chains of states for C&C channels, aggregated by 4-tuples (source/destination IP, port, protocol). Each chain was stored as MM, and detection involved checking if new flow chains matched pre-trained models. Their method showe effectiveness in distinguishing botnet traffic despite variability across C&C channels.
In [79], the authors present M&M model, a multilevel framework that captures the complex short- and long-term dynamics of 802.15.4 wireless links using HMMs and Mixtures of Multivariate Bernoullis (MMBs). This method models link behavior across time scales and generates synthetic traces that match real data in terms of packet reception rate statistics, run-length distributions, and conditional delivery patterns.
Limitations: The literature reviewed in this subsection highlights a fundamental trade-off inherent to Markov-based traffic models. While increasing state complexity enables richer temporal dynamics, it also raises concerns regarding parameter estimation, scalability, and practical deployment [121, 185]. Moreover, even advanced variants such as hierarchical MMPPs and nested HMMs remain bounded by finite memory, limiting their ability to reproduce true long-range dependence. These constraints indicate that Markov models approximate correlation effects rather than fully capturing them.
Summary: In summary, MMs provide a structured and analytically tractable framework for modeling temporal behavior in network traffic. The studies discussed demonstrate their effectiveness in capturing oscillatory patterns, protocol-level dynamics, and queueing behavior [121, 185]. Although they do not achieve full self-similarity, their balance between expressiveness and interpretability makes them particularly valuable for performance analysis, anomaly detection, and scenario-driven traffic synthesis.
III-D Autoregression (AR)
@@
An AR model is a type of time series model where the current value of a time series is expressed as a linear combination of its previous values (lags) and a random error term. An autoregressive model of order $p$, denoted as AR(p), is defined as:
@@
where, $X_{t}$ is the value of the time series at time $t$, $c$ is a constant term, $\phi_{1},\phi_{2},\ldots,\phi_{p}$ are the autoregressive coefficients, $p$ is the number of previous time steps (lags) considered in the model, and $\epsilon_{t}$ is a random error term (white noise) with mean 0 and constant variance $\sigma^{2}$. These models are widely used in time series analysis. AR Integrated Moving Average (ARIMA) models extend the AR model by considering the moving average and non-stationarity. ARFIMA extends ARIMA by introducing fractional differencing to account for long-memory dependence.
Shu et al. [155] investigated ARFIMA for accurate traffic prediction, emphasizing the need to capture both SRD and LRD. They note that ARFIMA, Markov modulated processes, Transform-Expand-Sample (TES), and scene-based models can all capture SRD and LRD, but the complexity of ARFIMA fitting has limited adoption. To address this, they propose a simplified fitting procedure, observing that the order (p, q) remains stable over long periods, enabling more efficient fitting. They further introduce an adapted prediction method that provides an upper probability bound for forecasts. Their results show that ARFIMA achieves higher accuracy than AR models in h-step forecasting while requiring fewer parameters.
ARIMA/GARCH [206] is a non-linear time series model combining linear ARIMA with Generalized Autoregressive Conditional Heteroskedasticity (GARCH) to predict network traffic. Accurate prediction requires capturing SRD, LRD, large-scale self-similarity, and small-scale multifractality. ARIMA captures only SRD, ARFIMA captures SRD and LRD but not multifractal behavior, while Multifractal Wavelet Models capture multifractality but fail in prediction. Network traffic is also non-stationary and non-linear, motivating ARIMA/GARCH, where the GARCH component models dynamic, time-varying variance to capture burstiness. The authors propose a parameter estimation procedure and an adaptive prediction scheme, demonstrating superior performance over ARFIMA in forecasting network traffic.
Limitations: The AR-based models reviewed reveal that improved memory modeling comes at the cost of increased complexity. As noted in [155], ARFIMA’s ability to capture both SRD and LRD is offset by challenging parameter estimation, motivating simplified fitting strategies. Similarly, ARIMA/GARCH models introduce additional parameters to model non-stationarity and burstiness [206], increasing computational overhead. These characteristics limit the practicality of AR-based models for large-scale synthetic trace generation, especially when real-time adaptability is required.
Summary: Overall, autoregressive models offer strong predictive capabilities for network traffic analysis. The progression from AR to ARFIMA and ARIMA/GARCH reflects a shift toward capturing increasingly complex dependence structures [155, 206]. As evidenced in the reviewed studies, their primary strength lies in forecasting and short-term traffic prediction rather than full traffic synthesis, positioning them as complementary tools within broader modeling pipelines.
III-E Fractional Brownian Motion (FBM)
FBM is defined as a centered Gaussian process $B_{H}(t)$, $t\geq 0$ with the covariance function,
@@ III-E Fractional Brownian Motion (FBM)
Here, $H\in(0,1)$ is the Hurst parameter or the Hurst index. FBM has been considered a suitable choice for modeling network traffic due to its inherent self-similarity, a property that mirrors modern network traffic. The authors in [135] study the accuracy of simulation for network traffic using FBM. The authors describe simulation methods for FBM based on its spectral representation and the property of stationary increments. The indicated methods were used for the simulation of self-similar traffic and the loading process of telecommunication networks. The simulated FBM is then used as an input process to a linear system. The goal is to construct a model that accurately approximates FBM while also considering the system’s response. While the Gaussian nature of FBM aids in analyzing queuing behavior, its limitations in capturing short-term correlations and rich scaling properties of real traffic have led researchers to explore multifractal models [121]. However, these models were also difficult to manage due to their analytical complexity.
Limitations: While FBM provides a mathematically elegant representation of self-similar traffic, its limitations become apparent when modeling realistic network behavior. As discussed in [121], FBM fails to capture short-term correlations and richer scaling behavior observed in empirical traffic traces. Although multifractal models address these deficiencies, their analytical complexity reduces their usability in practical modeling and simulation tasks.
Summary: In summary, FBM serves as a foundational reference model for understanding long-range dependence in network traffic. Its use in simulation and system-response analysis demonstrates its value for theoretical evaluation [135]. However, due to its limited expressiveness at fine time scales, FBM is best regarded as a conceptual benchmark rather than a standalone solution for comprehensive traffic generation.
III-F Other
III-F1 Hierarchical State-Machine-Based (HSM) Models
@@
These have been introduced as a more accurate approach to modeling control-plane traffic in cellular networks (e.g., 4G LTE and 5G). Traditional models, such as PPs, have proven inadequate due to their inability to capture the bursty nature and long-tailed distributions observed in control-plane traffic. J. Meng et al. [114] demonstrated that a two-level HSM structure can effectively model control-plane traffic. The authors show that this structure enables the capture of event dependencies among different types of control-plane events. Furthermore, by using a Semi-Markov process, the model can accurately represent the time spent by a User Equipment in each state, overcoming the limitations of the MM. Additionally, an adaptive clustering scheme is included to account for variations in traffic behavior based on factors such as device type and time of day. The proposed model is experimentally validated by comparing the synthetic traffic generated by the model against real-world traffic traces.
Limitations: Despite their demonstrated accuracy, HSM models introduce substantial modeling and deployment complexity. As shown in [114], defining hierarchical states, adaptive clusters, and semi-Markov timing requires detailed domain knowledge and extensive data preprocessing. This specialization limits the generalizability of HSMs beyond control-plane traffic scenarios.
Summary: In literature, HSM models represent a targeted solution for control-plane traffic generation in cellular networks. By explicitly modeling event dependencies and state durations, they address limitations of memoryless statistical models [114]. Their effectiveness underscores the importance of protocol-aware modeling for control-plane analysis.
III-F2 K-Nearest Neighbors (KNN)